OpenStudy (anonymous):

Find all the real solutions for the polynomial equation x^4 − 5x^3 − 30x^2 − 30x − 216 = 0. 1. x = −4, 9,−1 2. x = −4, 9 3. x = ±1, 6, 10 4. x = 1,−6, 10 5. x = −10, 6 6. None of these

OpenStudy (anonymous):

so let me make sure i understand the problem... FInd all the real solutions for the polynomial equation \[x^{4}-5x^{3}-30x^{2}-30x-216=0\] 1. x = -4, 9, -1 2. x = -4, 9 3. x = 1, -1, 6, 10 4. x = 1, -6, 10 5. x = -10, 6 6. None of these is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... the best thing here would be to pick a value in an answer, try it and eliminate answers based on the result... for instance, start with x = 1. check to see if plugging in 1 for x in the equation gives you a zero

OpenStudy (anonymous):

\[(1)^{4}-5(1)^{3}-30(1)^{3}-30(1)-216\] \[1-5-30-30-216\] does that = 0?

OpenStudy (anonymous):

No.... this is what I got for the solutions x = 9 x = − 4 x = − √6·I x = √6·I

OpenStudy (anonymous):

even better...so looking at just the real solutions there (there are two of them) which possible answer does that match?

OpenStudy (anonymous):

#2 but shouldn't it inclue - sqrt6*1?

OpenStudy (anonymous):

that's an i not a 1...those last 2 are imaginary pairs...

OpenStudy (anonymous):

yeah I meant to put i

OpenStudy (anonymous):

a fourth degree polynolmial has 4 solutions... in this case, 2 of them are real solutions and the other 2 are imaginary... all imaginary solutions come in conjugate pairs.

OpenStudy (anonymous):

\[i\sqrt{6}\]and \[-i\sqrt{6}\] are not real solutions which is what was asked...so they wouldn't be in the list

OpenStudy (anonymous):

okay thanks